12. A reaction has Ea = 50.0 kJ mol G1. If the specific rate constant (k)
for this reaction is 0.00365 s G1 at 25 o C, what will the value of the
specific rate constant be at 100EC ?
Answer #1

Given:

T1 = 25 oC

=(25+273)K

= 298 K

T2 = 100 oC

=(100+273)K

= 373 K

K1 = 3.65*10^-3 s-1

Ea = 50 KJ/mol

= 50000 J/mol

use:

ln(K2/K1) = (Ea/R)*(1/T1 - 1/T2)

ln(K2/3.65*10^-3) = (50000.0/8.314)*(1/298 - 1/373.0)

ln(K2/3.65*10^-3) = 6014*(6.747*10^-4)

ln(K2/3.65*10^-3) = 4.058

(K2/3.65*10^-3) = e^(4.058)

(K2/3.65*10^-3) = 57.85

K2 = 0.2112 s-1

Answer: 0.211 s-1